# Vision-Language Models for Edge Networks: A Comprehensive Survey

### v1

@@ I Introduction
Figure 1: An Overview of The Role of VLMs in IoT Applications.
@@ I-B Market Statistics and Research Trends
VLMs have rapidly emerged as a new market in recent years, fueled by the rising demand for intelligent systems that can understand and reason with both visual and textual information. The global AI market was valued at USD 58.3 billion in 2021, and it is expected to reach USD 309.6 billion by 2026, with a CAGR of 39.7% [42]. This segment of the VLM market is projected to grow at the most rapid rate. The overall VLM market is anticipated to grow significantly over the projection period. The global market size of VLMs is estimated to reach around \$2.5 billion in 2024, increasing from \$1.8 billion in 2023 and \$1.2 billion in 2022 [43]. Why is everyone talking about it? Because VLMs find application in diverse sectors, including healthcare, automotive, and consumer electronics. For example, one of the factors driving the growth of the VLMs market is the adoption of VLMs to develop ADAS and autonomous driving solutions in the automotive industry. At the same time, with the increase in smart devices and the IoT, the need for lightweight VLMs that perform well on edge devices has become increasingly urgent.
@@
Research trends in VLMs indicate a strong focus on enhancing model efficiency and accuracy while reducing computational overhead. Recent studies have explored various techniques for model compression, including pruning, quantization, and knowledge distillation [9, 11, 10]. Additionally, there is growing interest in developing new architectures that leverage the strengths of both CNNs and transformer models [26, 29]. These hybrid models aim to balance the computational efficiency of CNNs with the powerful representation capabilities of transformers. Another emerging trend is using multi-task learning frameworks, where a single VLM is trained to perform multiple related tasks, improving overall efficiency and reducing the need for task-specific models. Notably, the number of research papers published on VLMs and AI on edge devices has increased significantly, reflecting the growing academic interest in this field.
@@
The applications of VLMs are expanding rapidly, with significant investments being made in sectors such as healthcare, retail, and security. In healthcare, VLMs are utilized for tasks such as medical image analysis, disease diagnosis, and telemedicine, providing real-time assistance to healthcare professionals [30, 31]. The retail sector is leveraging VLMs for enhanced customer experiences through smart shopping assistants and personalized marketing [32, 33]. Security applications include automated surveillance systems that can analyze and interpret visual data to detect anomalies and potential threats. These applications demonstrate the versatility and impact of VLMs across various industries, driving further research and development in this field. AI development on edge devices has become a critical area of focus due to the need for real-time processing, reduced latency, and improved privacy. Edge AI involves deploying AI models directly on devices such as smartphones, cameras, and IoT sensors, enabling local data processing without relying on cloud infrastructure [17]. This shift towards edge computing is driven by the limitations of cloud-based AI, including latency issues, bandwidth constraints, and data privacy concerns. Research in edge AI is focused on optimizing model architectures and developing specialized hardware accelerators to support efficient inference on resource-constrained devices [12, 13]. Companies like NVIDIA, Intel, and Google invest heavily in edge AI solutions, indicating a robust market growth trajectory. According to Allied Market Research, the global edge AI hardware market is expected to reach USD 3.89 billion by 2025, growing at a CAGR of 20.6% from 2018 to 2025 [44].
@@ I-C Existing Surveys and Tutorials
Few surveys and tutorials have reviewed VLMs, their efficiency, and applications [34, 35, 36, 37, 38, 39, 40, 41]. Table I summarizes some of this work Scopes and how it is different than ours. The authors in[34] focused on vision-language pre-trained models, discussing the evolution of these models, different architectures used, and methods for integrating vision and language modalities. Another work [35] explored vision-language intelligence, emphasizing tasks, representation learning, and the development of large models. They provided insights into the performance improvements and future research directions in this area. Xing et al. [36] surveyed efficient fine-tuning methods for vision-language models, focusing on Prompt and Adapter techniques. They discussed various strategies to enhance fine-tuning efficiency and addressed challenges related to efficient fine-tuning. Ghosh et al. [37] provided a comprehensive overview of the current methodologies and future directions of vision-language models, highlighting the strengths and limitations of existing approaches and suggesting areas for further exploration. Zhang et al. [38] surveyed vision-language models for vision tasks, discussing the theoretical foundations, practical applications, and identifying challenges and opportunities in applying these models in fields like medical imaging and industrial automation. Another survey [39] focused on multimodal large language models for autonomous driving, discussing the integration of different modalities, methodologies to enhance model performance, and specific applications in autonomous driving scenarios. Yin et al. [40] surveyed multimodal large language models with a focus on efficient design and diverse applications, covering architectures, strategies to enhance efficiency, and applications in fields like biomedical analysis and document understanding. Lastly, Jin et al. [41] provided a survey on efficient multimodal large language models, discussing methods to reduce computational costs, improve efficiency, and applications in areas like high-resolution image understanding and medical question-answering, highlighting future research directions and challenges in the field.
@@ II-A Key Concepts
Assuming a text input sequence $T=[t_{1},t_{2},\ldots,t_{N}]$ where each $t_{i}$ is the $i$-th token, the representation for $T$ can be obtained through tokenization and embedding. Steps in the Flow: Each token $t_{i}$ is mapped to a high-dimensional word embedding $\mathbf{e}_{i}$ that captures the semantic features of the word. This mapping is provided by a pretrained embedding function (often the output of transformer-based models like BERT, or other pretrained language models [45, 46]):
$\mathbf{e}_{i}=\text{Embedding}(t_{i}).$ (1)
In order to account for this sequence information, we simply add positional embeddings $\mathbf{p}_{i}$ to each of the token embeddings. These positional embeddings give a sense of word position, which is essential for maintaining the structure of the text in tasks that rely on understanding word relationships [47]:
$\mathbf{h}_{i}=\mathbf{e}_{i}+\mathbf{p}_{i}.$ (2)
This word + positional embedding $\mathbf{h}_{i}$ is then forwarded through a number of transformer layers to allow the model to assemble a rich contextual representation of the text input. These contextualized embeddings are used to derive a single representation that joins together with the visual features.
@@
For the image input $I$, the representation appropriately consists of obtaining informative visual features that inform the image’s semantic and spatial content. This approach is most often implemented with either convolutional neural networks (CNNs) or Vision Transformers (ViTs), depending on the model’s architecture. The image $I$ is generally decomposed into a grid of patches, where each patch $I_{j}$ represents a local image region. The next step consists of encoding each patch into a feature vector $\mathbf{v}_{j}$ representing its visual content [48, 49]:
$\mathbf{v}_{j}=\text{VisionEncoder}(I_{j}).$ (3)
The feature vectors $\mathbf{v}_{j}$ are obtained from the last convolutional layers for CNN-based models. In contrast, for ViTs, each image patch is linearly embedded and further processed using self-attention layers. This yields a set of visual embeddings containing both low-level and high-level features. The embeddings are mapped with text embeddings in a multimodal representation space to enable cross-modal tasks.
@@ II-B Mechanisms for Vision-Language Interaction
In contrast, single-stream models do early fusion by interleaving visual and textual encodings into a single sequence fed through a common encoder — often, a transformer [50, 51]. This architecture relies on the assumption that a single transformer encoder can adequately model the interactions among the modalities. This means that the language and image tokens are tokenized and embedded and then combined into one sequence — the model processes them together, having the ability to learn visual and textual attributes at the same time. This approach encodes the two modalities using this common representation, which can help efficiently model the intricate relationships and interactions between them. In the single-stream framework, text embeddings $\mathbf{h}_{i}$ and image embeddings $\mathbf{v}_{j}$ are concatenated and processed through a transformer [50, 51]:
$\mathbf{z}_{k}=\text{Transformer}([\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{N};\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}]).$ (4)
@@
II-C Efficient Fine-Tuning Methods for Vision-Language Models
II-C1 Fine-tuning with Prompts
II-C2 Adapter-Based Fine-Tuning
II-D Existing VLM Models
Model Year Fusion Scheme Parameters Applications
MobileVLM [68] 2024 Single stream 1.1B Mobile applications, Real-time image captioning
LightCLIP [69] 2024 Dual stream 2.45M Image classification, Zero-shot learning
MoE-TinyMed [70] 2024 Single stream Not specified Medical imaging, Diagnostic assistance
PaLI-3 [71] 2024 Single stream Not specified Image captioning, Object detection
RegionGPT [72] 2024 Dual stream Not specified Regional image analysis, Multimodal translation
Ins-DetCLIP [73] 2024 Single stream Not specified Object detection, Scene understanding
CLIP-Adapter [67] 2023 Dual stream Not specified Image-text retrieval, Few-shot learning
Lightweight Unsupervised Federated Learning [78] 2023 Dual stream Not specified Distributed learning, Privacy-preserving training
CLIP [80] 2022 Dual stream 400M Zero-shot learning, Image-text matching
VisualBERT [81] 2022 Single stream 110M Visual question answering, Image captioning
@@ II-D Existing VLM Models
Figure 4: The MobileVLM architecture. Inputs include visual data $X_{v}$ and textual queries $X_{q}$, processed by a vision encoder and tokenizer, respectively, producing hidden states $H_{v}$ and $H_{q}$. These are fed into MobileLLaMA, which outputs a response $Y_{a}$. The Lightweight Downsample Projector (LDP) efficiently processes the visual input using depthwise and pointwise convolutions [68].
@@ III-A Existing Low Complexity VLMs
Many papers and models have been developed to run VLM on edge devices. EdgeVL is a novel framework designed to adapt large VLMs for edge devices by leveraging dual-modality knowledge distillation and quantization-aware contrastive learning. Fig.5 illustrates how the model focuses on efficiently aligning features from RGB and non-RGB images without manual annotation, making it versatile for various visual modalities. EdgeVL achieves up to a 15.4% improvement in accuracy and a 93-fold reduction in model size, which is crucial for deployment on resource-limited devices. The model’s design allows for a streamlined adaptation process, where a student encoder is trained to mimic a large, pre-trained teacher model like CLIP, ensuring high-quality feature extraction despite the reduced model size [88].
@@ III-B Deployment of VLMs on Edge Devices
In order to deploy the VLM model on edge devices, several important steps are required to achieve efficient deployment. These steps (as shown in Fig. 6) include:

### v2

@@ I Introduction
Figure 1: Overview of integrating Vision-Language Models (VLMs) into edge networks. The diagram highlights IoT applications (e.g., healthcare, autonomous driving, smart homes, surveillance, gaming, education, fitness tracking, Industry 4.0, sports) that benefit from VLMs to address traditional Machine Learning (ML) challenges such as limited training data and semantic reasoning. It outlines key ML techniques (transfer learning, federated learning, semi-supervised learning) and model optimization methods (knowledge distillation, pruning, compression) for efficient deployment. Crucial enabling technologies, including edge computing, quantum communication and computing, and advanced wireless networks (5G/6G), are also presented to support VLM deployment at the edge
@@ I-B Market Statistics and Research Trends
VLMs have rapidly emerged as a new market, driven by demand for systems that understand and reason with visual and textual information. The global AI market was valued at USD 58.3 billion in 2021 and is projected to reach USD 309.6 billion by 2026, with a CAGR of 39.7% [42]. The VLM segment is expected to grow at the fastest rate. Its market size is estimated to reach \$2.5 billion in 2024, up from \$1.8 billion in 2023 and \$1.2 billion in 2022 [43]. VLMs are gaining traction across sectors such as healthcare, automotive, and consumer electronics. In the automotive industry, VLMs contribute to ADAS and autonomous driving solutions. With the rise of smart devices and IoT, there is increasing demand for lightweight VLMs capable of running efficiently on edge devices.
@@
Research trends in VLMs emphasize improving efficiency and accuracy while minimizing computational cost. Studies have explored model compression techniques such as pruning, quantization, and knowledge distillation [9, 11, 10]. There is also growing interest in hybrid architectures combining CNNs and transformers [26, 29], aiming to merge CNNs’ efficiency with transformers’ representational power. Multi-task learning is another trend, where a single VLM handles multiple tasks, enhancing efficiency and reducing reliance on task-specific models. The number of research papers on VLMs and AI for edge devices has notably increased, reflecting heightened academic interest.
@@
VLM applications are expanding rapidly, with major investments in healthcare, retail, and security. In healthcare, they support medical image analysis, diagnosis, and telemedicine [30, 31]. Retail uses include smart shopping assistants and personalized marketing [32, 33]. In security, automated surveillance leverages VLMs to detect anomalies and threats. These examples show VLMs’ versatility and growing impact. Edge AI, crucial for real-time processing and privacy, deploys models on devices like smartphones and IoT sensors, bypassing cloud dependency [17]. This shift is motivated by latency, bandwidth, and privacy issues. Research in edge AI focuses on model optimization and hardware accelerators for efficient inference on constrained devices [12, 13]. Companies like NVIDIA, Intel, and Google are investing heavily in this area. The global edge AI hardware market is projected to reach USD 3.89 billion by 2025, growing at a CAGR of 20.6% from 2018 [44].
@@ I-C Existing Surveys and Tutorials
Few surveys have reviewed VLMs, their efficiency, and applications [34, 35, 36, 37, 38, 39, 40, 41]. Table I summarizes key scopes and how they differ from ours.
[34] focused on vision-language pre-trained models, covering their evolution, architectures, and integration methods. [35] explored vision-language intelligence, emphasizing tasks, representation learning, and large model development, along with performance insights and research directions. Xing et al. [36] surveyed efficient fine-tuning methods for VLMs, focusing on Prompt and Adapter techniques and related challenges. Ghosh et al. [37] provided an overview of current methodologies and future directions, highlighting strengths, limitations, and areas for exploration. Zhang et al. [38] discussed VLMs for vision tasks, covering theoretical foundations, practical applications, and challenges in domains like medical imaging and industrial automation. [39] focused on multimodal models for autonomous driving, discussing modality integration, performance methods, and specific applications. Yin et al. [40] reviewed multimodal large language models with emphasis on efficient design, architectures, and applications in biomedical analysis and document understanding. Finally, Jin et al. [41] surveyed efficient multimodal LLMs, focusing on reducing computational cost, improving efficiency, and applications in high-res image understanding and medical QA, while outlining challenges and research directions.
@@ II-A Key Concepts
Assuming a text input sequence
$T=[t_{1},t_{2},\ldots,t_{N}],$
where each $t_{i}$ is the $i$-th token (e.g., an integer ID or a subword unit). We first map each token $t_{i}$ into a high-dimensional word embedding $\mathbf{e}_{i}\in\mathbb{R}^{d}$. Concretely, one may define a learnable embedding matrix $\mathbf{W}_{E}\in\mathbb{R}^{d\times|\mathcal{V}|}$ (where $|\mathcal{V}|$ is the vocabulary size), so that each $t_{i}$ (represented as a one-hot or index) is projected into $\mathbf{e}_{i}$:
$\mathbf{e}_{i}=\mathbf{W}_{E}\cdot\mathrm{one\_hot}(t_{i})+\mathbf{b}_{E},$ (1)
where $\mathbf{b}_{E}\in\mathbb{R}^{d}$ is a bias term and $\mathrm{one\_hot}(t_{i})\in\mathbb{R}^{|\mathcal{V}|}$ is a sparse vector with 1 at position $t_{i}$ and 0 elsewhere. Equivalently, modern frameworks often bypass the explicit one-hot representation by using an embedding lookup.
In order to account for the sequence information, we add a positional embedding $\mathbf{p}_{i}\in\mathbb{R}^{d}$ to each token embedding. Often, $\mathbf{p}_{i}$ is either a learned parameter or derived from a fixed function (e.g., sinusoidal) [45]. The total embedding $\mathbf{h}_{i}$ of the $i$-th token is then:
$\mathbf{h}_{i}=\mathbf{e}_{i}+\mathbf{p}_{i},$ (2)
where $\mathbf{h}_{i}$ denotes the final representation of the $i^{th}$ token. $\mathbf{e}_{i}$ denotes the embedding vector. Moreover, $\mathbf{p}_{i}$ represents the positional encoding. These combined embeddings $\{\mathbf{h}_{i}\}_{i=1}^{N}$ are subsequently processed through multiple transformer layers or other neural encoders to build contextualized text representations. For example, denoting one transformer block by $\mathrm{TransBlock}(\cdot)$, a simplified representation of one layer’s output $\mathbf{h}_{i}^{(l)}$ (for the $l$-th layer) could be:
$\mathbf{h}_{i}^{(l)}=\mathrm{TransBlock}\bigl{(}\{\mathbf{h}_{1}^{(l-1)},\ldots,\mathbf{h}_{N}^{(l-1)}\}\bigr{)}_{i},$ (3)
where $\mathbf{h}_{i}^{(0)}=\mathbf{h}_{i}$. Stacking multiple layers refines the representation of each token based on its context in the sequence. Ultimately, one can use $\{\mathbf{h}_{i}^{(L)}\}$ from the final layer $L$ for downstream tasks or for multimodal fusion.
@@
For the image input $I$, we extract informative visual features that reflect the image’s semantic and spatial content. Depending on the model’s backbone, either convolutional neural networks (CNNs) or Vision Transformers (ViTs) can be used. Let us divide the image $I$ into patches $\{I_{j}\}_{j=1}^{M}$, where each patch covers a local region of the image. A learnable vision encoder is then applied to each patch:
$\mathbf{v}_{j}=\mathrm{VisionEncoder}(I_{j}),$ (4)
with $\mathbf{v}_{j}\in\mathbb{R}^{d}$. In CNN-based backbones, $\mathrm{VisionEncoder}(\cdot)$ might represent the final convolutional layers, extracting feature maps that encode both low-level and high-level cues. In Vision Transformers [46], one often computes a linear projection for each patch, adds learnable 2D positional embeddings, and processes them through multi-head self-attention blocks, analogous to the transformer encoder in text processing. Concretely, for a ViT, one may write
$\mathbf{v}_{j}^{(0)}=\mathrm{PatchEmbed}(I_{j})+\mathbf{p}_{j}^{\text{(img)}},$
followed by repeated transformer layers:
$\mathbf{v}_{j}^{(l)}=\mathrm{TransBlock}\bigl{(}\{\mathbf{v}_{1}^{(l-1)},\ldots,\mathbf{v}_{M}^{(l-1)}\}\bigr{)}_{j},$ (5)
to yield the final visual tokens $\{\mathbf{v}_{j}^{(L)}\}$.
The result $\{\mathbf{v}_{j}\}_{j=1}^{M}$ (or the final-layer embeddings in the case of a ViT) provides a set of patch-level representations that can be mapped into the same dimensional space as $\mathbf{h}_{i}$. By aligning $\mathbf{v}_{j}$ and $\mathbf{h}_{i}$ in a unified embedding space, the model can learn cross-modal correspondences.
@@ II-B Mechanisms for Vision-Language Interaction
In contrast, single-stream models do early fusion by interleaving visual and textual encodings into a single sequence fed through a common encoder — often, a transformer [47, 48]. This architecture relies on the assumption that a single transformer encoder can adequately model the interactions among the modalities. Concretely, denote the final text embeddings by $\{\mathbf{h}_{i}^{(L_{t})}\}_{i=1}^{N}$ and the final image embeddings by $\{\mathbf{v}_{j}^{(L_{v})}\}_{j=1}^{M}$. These are concatenated into a unified sequence:
$[\mathbf{h}_{1}^{(L_{t})},\mathbf{h}_{2}^{(L_{t})},\ldots,\mathbf{h}_{N}^{(L_{t})};\ \mathbf{v}_{1}^{(L_{v})},\mathbf{v}_{2}^{(L_{v})},\ldots,\mathbf{v}_{M}^{(L_{v})}].$
A shared multimodal transformer, $\mathrm{Transformer}(\cdot)$, is then applied to capture deep interactions between the textual and visual tokens. Formally, the output embedding $\mathbf{z}_{k}$ at position $k$ in the fused sequence is:
$\mathbf{z}_{k}=\mathrm{Transformer}\Bigl{(}\bigl{[}\mathbf{h}_{1}^{(L_{t})},\ldots,\mathbf{h}_{N}^{(L_{t})};\ \mathbf{v}_{1}^{(L_{v})},\ldots,\mathbf{v}_{M}^{(L_{v})}\bigr{]}\Bigr{)}_{k}.$ (6)
This single-stream approach encodes both modalities together, allowing each token (text or patch) to directly attend to and influence every other token. Such early fusion can be advantageous for tasks demanding fine-grained alignment between language and vision, such as Visual Question Answering and grounded dialogue systems.
@@
Model Year Fusion Scheme Parameters Applications
MobileVLM V2 [50] 2024 Single stream 1.1B Mobile applications, Real-time image captioning
ViTamin [55] 2024 Single stream Not specified Image classification, Open-vocabulary detection
LINGO-2 [56] 2024 Single stream 5B Autonomous driving, Driving behavior prediction
InstructBLIP [57] 2024 Single stream Not specified Instruction tuning, Question answering, Image captioning
RAVEN [58] 2024 Single stream Not specified Visual question answering, Image captioning
ScreenAI [59] 2024 Single stream Not specified UI understanding, Infographic analysis
@@
II-C Existing Lightweight VLM Models
@@
EfficientVLM is a vision-language model designed for efficiency, utilizing knowledge distillation and modal-adaptive pruning to reduce size and enhance speed. With 93 million parameters, it achieves 98.4% of its teacher model’s performance while accelerating inference by 2.2 times, demonstrating strong results in tasks like visual question answering and image retrieval [8].
Unified-IO is a versatile model capable of handling a wide array of tasks across vision, language, and multimodal domains. By converting diverse inputs and outputs into a unified sequence of discrete tokens, it employs a single transformer-based architecture trained on over 90 datasets. This approach enables Unified-IO to perform tasks such as pose estimation, object detection, and question answering without task-specific fine-tuning [51].
MiniVLM is a compact vision-language model focused on efficiency. It integrates a Two-stage Efficient feature Extractor (TEE) and a transformer-based fusion module, resulting in a 73% reduction in model size and a 94% decrease in inference time compared to larger counterparts. Despite its smaller size, MiniVLM retains 94-97% of the accuracy on various vision-language tasks, making it suitable for edge applications [61].
@@
Figure 4: The MobileVLM architecture. Inputs include visual data $X_{v}\in\mathbb{R}^{N_{v}\times D_{v}}$ and textual queries $X_{q}$, where $N_{v}$ is the number of visual tokens and $D_{v}$ is the visual feature dimension. These are processed by a vision encoder and tokenizer, respectively, producing hidden states $H_{v}\in\mathbb{R}^{(N_{v}/4)\times D_{v}}$ and $H_{q}$. The visual features are passed through a Lightweight Downsample Projector (LDP), which efficiently compresses the input using depthwise and pointwise convolutions. The resulting features $H_{v}$ and $H_{q}$ are then fed into MobileLLaMA, a compact vision-language model, which generates the final response $Y_{a}$.[50].
II-D Efficient Fine-Tuning Methods for Vision-Language Models
II-D1 Fine-tuning with Prompts
II-D2 Adapter-Based Fine-Tuning
@@ III-A Existing Low Complexity VLMs
While all edge-deployed Vision-Language Models (VLMs) must be lightweight, not all lightweight VLMs are specifically optimized for edge deployment. Lightweight VLMs focus on reducing model size, computational cost, and latency while maintaining strong performance, but they may still rely on cloud-based processing. In contrast, edge-specific VLMs are designed with additional constraints such as energy efficiency, low-latency operation, and privacy-preserving inference, making them more suitable for real-time applications on resource-constrained devices.
Several models have been developed for edge environments, with varying degrees of optimization. One such example is EdgeVL, a framework designed to adapt large-scale models for edge deployment. Unlike general lightweight VLMs, which primarily aim to reduce size and complexity, EdgeVL incorporates dual-modality knowledge distillation and quantization-aware contrastive learning to align representations between a large teacher model, such as CLIP, and a compact student model. This enables efficient processing of RGB and non-RGB images without manual annotation, improving adaptability across vision tasks. The framework achieves significant improvements in accuracy and a substantial reduction in model size, making it highly suitable for real-world edge applications [87].
Other models, such as MiniVLM and LightVLP, represent general-purpose lightweight VLMs that, while optimized for efficiency, are not explicitly designed for edge deployment. MiniVLM reduces visual feature extraction time by 95% compared to baseline models and achieves competitive performance on various downstream tasks [88]. LightVLP employs a gated interaction mechanism to handle noise in aligned image-text pairs, resulting in efficient and effective vision-language pre-training [89]. However, these models do not incorporate edge-specific techniques such as on-device inference, federated learning, or adaptive computation, making them more suited for cloud-assisted environments or mobile applications rather than fully autonomous edge computing.
Selecting the appropriate model depends on deployment constraints. General lightweight models may be effective in low-power cloud-assisted settings, while edge-specific models integrate optimizations that enable local inference with minimal latency and energy consumption, making them essential for IoT, autonomous systems, and privacy-sensitive applications.
@@ III-B Deployment of VLMs on Edge Devices
Deploying VLMs on edge devices requires several key steps for efficiency, as illustrated in Fig. 6: